# Birgir Hrafnkelsson Morris Riedel Egill Skúlason

### v1

@@ 3.1.1 Kohn–Sham DFT: exact in principle, orbital constrained by construction
The Hohenberg–Kohn [56] theorem formulates the ground-state energy in terms of the electron density. Kohn-Sham (KS) density functional theory (DFT) builds a noninteracting system of virtual electrons that reproduces the interacting density, which in turn brings the conpt of orbitals back to the fore. The KS equations employ a single multiplicative potential
@@
If $E_{xc}[n]$ were known, we could recover the exact density and energy. In practice we choose approximations [57] (LDA/GGA/meta-GGA) and gain correlation at modest cost. This local-potential form dovetails with radial finite elements and enables a fast, stable self-consistent loop.
The Hartree potential is found by solving the Poisson equation, $\nabla^{2}V_{H}({\bf x})=-4\pi n({\bf x})$, where $n(r)$ is the radial electron density constructed from the wave functions. Because the potential depends on the wave functions and the wave functions depend on the potential, these equations must again be solved iteratively until self-consistency is achieved.
@@ 3.2 The Physical and Mathematical Problem
The electronic structure of an isolated, spherically symmetric atom can be described at two primary levels of theory, depending on the level of rigor, the non-relativistic Schrödinger equation and the relativistic Dirac equation.
@@
Since the KS framework of DFT maps the complex many-body problem onto a tractable set of single-particle equations, a cyclic depence arises, where the potential depends on the wave functions (via the electron density) and the wave functions depend on the potential. The KS equations must be solved iteratively until self constency is achieved. The governing mathematical models for the WaveFunction entity are thus the radial Schrödinger and Dirac equations, solved within this self-consistent loop  [^6] In passing we note that larger molecules and complexes are treated often with non-relativistic methods since core electrons are represented with a pseudo-potential or projected augmented wave [58].
@@ 3.2.2 The Radial Dirac Equation
For heavy atoms, the appropriate single-particle theory is relativistic. The central-field Dirac equation leads to two coupled first-order radial equations for the large and small components $(P_{n\kappa},Q_{n\kappa})$ of a four-component spinor, with the relativistic quantum number $\kappa$ encoding $(l,j)$ [59].
@@
3.3 A Multi-Layered Representational Strategy
The featom code employs a high-order finite element method to solve the governing equations. The radial coordinate is discretized into a mesh of finite elements, and within each element, the solution is expanded in a basis of high-order Lagrange polynomials defined on Gauss-Lobatto nodes. This spectral element approach yields exponential convergence with respect to the polynomial order, providing high accuracy with a relatively small number of basis functions. The success of this approach, however, rests on a cascade of intelligent choices in representation at the mathematical, numerical, and software levels.
@@ 3.3.1 Layer 1: The Mathematical Representation (Squared Hamiltonian)
Direct discretization of the Dirac Hamiltonian operator is notoriously difficult because its energy spectrum is unbounded from both above and below, leading to spurious, unphysical solutions. To circumvent this, a different mathematical representation is used: we solve the eigenvalue problem for the square of the Dirac Hamiltonian, $(H+\mathbb{I}c^{2})^{2}$. This squared operator has the same eigenfunctions as the original operator, and its eigenvalues are simply the square of the original eigenvalues, $(E+c^{2})^{2}$. Crucially, the squared operator is bounded from below, making it amenable to standard variational techniques like the finite element method without generating spurious states.
@@ 3.3.4 Layer 4: The Software Representation (Modern, Maintainable Code)
This robust code is supported by a professional software engineering infrastructure. This work involved introducing the flexible Meson build system alongside the existing Fortran Package Manager (fpm), establishing a comprehensive automated test harness, and refining continuous integration (CI) pipelines. This focus on the software representation ensures correctness through automated validation, lowers the barrier for collaboration, and guarantees long-term maintainability and scientific reproducibility.
@@ 3.4 From KS to HF: conceptually simple, practically subtle in spherical FE
A simple extension towards an HF/KS implementation in the radial FE code follows a “local-in-the-loop, exact-after” strategy 777gh-26 to featom:
@@ 3.5 Performance and accuracy
As designed, the featom code is tailored towards being state-of-the-art as a DFT solver for relativistic calculations, which are otherwise treated only through expensive quantum chemical approaches [66]. Figure 9 demonstrates the systematic, reproducible convergence of the solver for uranium in three complementary regimes: domain truncation, \$p\$- and \$h\$-refinement, and the achieved energy precision.
@@
The efficiency of the implementation, measured by wall time, was benchmarked against the state-of-the-art shooting-method code dftatom. For a DFT calculation of uranium converged to an accuracy of $10^{-6}$ Hartree, featom shows a significant speedup for non-relativistic calculations and competitive performance for relativistic ones, validating the effectiveness of the chosen representations.
@@
The finite element approach implemented in featom provides a state of the art performance for relativistic atomic structure calculations. The work on this tool exemplifies the central thesis that overcoming computational barriers in science requires a holistic approach to representation. The accuracy of the physics is enabled by successively more effective representations, namely, a mathematically stable squared Hamiltonian, a functionally smooth set of corrected wave functions, and a numerically robust algorithm for their integration. This entire structure is captured in a final, crucial representation, which is the software itself, using modern design and robust engineering practices. This concrete artifact ultimately makes the abstract theories physically relevant by enabling their efficient and reliable computation. However, the converse is also true, not every “conceptually simple” mapping (e.g. Hartree-Fock $\to$ “a potential”) respects the representation. Where the physics demands an operator (HF exchange), we either keep the operator or solve a re-representation problem, e.g. through an optimized effective potential [65, 67, 68]. We will return to this concept in later chapters; when we formulate efficient reaction-path searches and have Gaussian Process accelerators succeed because their internal representations (forces, curvatures, kernels) are chosen to make the numerics stable and the computation scalable.
@@ 4.1 Bonding analysis
To unambiguously distinguish covalently bonded molecular fragments from transient non-covalent contacts, a robust analysis of the system’s bonding network is essential. This can be approached from two distinct perspectives, namely a heuristic geometric definition or a more rigorous definition based on the system’s electronic structure. We implement in rgpycrumbs both methods, with a pyvista backend 888inspired by solvis allowing for a flexible and chemically aware analysis of molecular connectivity.
@@ 4.1.1 Geometric method: Covalent cutoff
The simplest and most computationally efficient method for defining a bonding network is based on geometry. In this approach, we define that a bond exists between two atoms, $i$ and $j$, based on their interatomic distance, $d_{ij}$. Specicially, the bond exists when atoms are closer to each other lthan a scaled sum of their tabulated covalent radii, $r_{i}$ and $r_{j}$. This relationship is governed by the inequality:
@@
Here, $M$ is a dimensionless scaling multiplier 999typically between $1.1$ and $1.3$ used to adjust the strictness of the criterion. While this method is extremely fast, and widely available due to being the method used by the ASE [74] graphical user interface (GUI); a purely geometric measure for molecules is a significant drawback, as the base unit of calculations are centered on electrons (Chapter 3). Such measures lack “chemical intuition” and can fail in sterically crowded environments where non-bonded atoms are forced into close proximity, leading to the false identification of covalent bonds (as illustrated in Figure 10 B and C.
@@ 4.1.2 Electronic density: Wiberg bond order
Such a measure can be significantly more robust than the geometric approach as it hinges on the calculation of the actual chemical interactions. It can reliably distinguish between genuine covalent bonds, which have significant shared electron density (typically WBO > 0.7), and close non-covalent contacts, which exhibit negligible bond orders. As shown in Figure 10, the WBO between sterically close but non-bonded atoms evaluates to a near zero value, correctly identifying them as belonging to separate molecular fragments. The primary trade-off for this increased accuracy is the higher computational cost associated with performing the underlying electronic structure calculation, which is largely alleviated by using GFN2-xTB semi-empirical calculation [76] 101010this is still not quick enough for extended systems however..
@@
4.2 Path visualization
@@
To visualize the local structure of the potential energy surface, we interpolate the scattered energies onto the RMSD plane. This is achieved using a Clough-Tocher 2D interpolator [80, 81], which is the method implemented by the griddata function with the cubic option in SciPy [77].
This method first constructs a Delaunay triangulation $\mathcal{T}$ of the set of points $\{(x_{i},y_{i})\}$ [82, 83]. The interpolant $E(x,y)$ is then defined as a piecewise bivariate cubic polynomial. For any point $(x,y)$ that falls within a specific triangle $\Delta_{k}\in\mathcal{T}$, the energy is given by a unique cubic polynomial $C_{k}(x,y)$:
$E(x,y)=C_{k}(x,y)=\sum_{i+j\leq 3}a_{ij}^{(k)}x^{i}y^{j}\quad\text{for}\quad(x,y)\in\Delta_{k}\in\mathcal{T}$ (53)
The coefficients $a_{ij}^{(k)}$ for each triangle are determined not only by the energy values $E_{i}$ at the triangle’s vertices but also by enforcing C1 continuity (continuous first derivatives) across the boundaries with adjacent triangles. This C1 continuity guarantees that the resulting interpolated surface is smooth and its gradient, $\nabla E$, is well-defined everywhere, which is physically necessary for representing a continuous potential energy surface. This produces the smooth contour maps used to visualize the topography of the energy landscape, shown in section 4.3.3, Figure 12.
@@ 4.3 EON
Most of the calculations in this thesis go through EON. Rather than exessively modifying the SVN copy, a new release was drafted, v2.8.0 121212accompanying documentation : https://eondocs.org. All the methods presented in this thesis are either the direct result of, or stemmed from the landmark modernization of the EON client code. This multi-year development effort, spawning millions of lines of code and documentation, overhauled the entire framework to be more powerful, flexible, and robust, transforming it from a legacy tool into a modern scientific platform. The effort focused on several key areas of software engineering.
@@ 4.3.1 Eliminating I/O Bottlenecks with a Client-Server Architecture
This effort required not only developing the server architecture within EON 131313gh-244 to EON but also contributing directly to the NWChem codebase to improve its capabilities as a client. A key pull request, which was merged into the official NWChem repository 141414gh-1145 to NWChem , enhanced its socket client with a robust polling and retry mechanism. This modification allows the NWChem client to wait patiently for the EON server to become available, a critical feature for ensuring stable, loosely-coupled communication between the two persistent programs.

### v2

@@ 3.1.1 Kohn–Sham DFT: exact in principle, orbital constrained by construction
The Hohenberg–Kohn [69] theorem formulates the ground-state energy in terms of the electron density. Kohn-Sham (KS) density functional theory (DFT) builds a noninteracting system of virtual electrons that reproduces the interacting density, which brings the concept of orbitals back to the fore. The KS equations employ a single multiplicative potential
@@
If $E_{xc}[n]$ were known, we could recover the exact density and energy. In practice we choose approximations [70] (LDA/GGA/meta-GGA) and gain correlation corrections at modest cost. This local-potential form dovetails with radial finite elements and enables a fast, stable self-consistent loop.
We find the Hartree potential by solving the Poisson equation, $\nabla^{2}V_{H}({\bf x})=-4\pi n({\bf x})$, where $n(r)$ represents the radial electron density constructed from the wave functions. Because the potential depends on the wave functions and the wave functions depend on the potential, these equations require iterative solution until they reach self-consistency.
@@ 3.2 The Physical and Mathematical Problem
We describe the electronic structure of an isolated, spherically symmetric atom at two primary levels of theory, depending on the required rigor: the non-relativistic Schrödinger equation and the relativistic Dirac equation.
@@
Since the KS framework of DFT maps the complex many-electron problem onto a tractable set of single-particle equations, a cyclic dependence arises: the potential depends on the wave functions (via the electron density) and the wave functions depend on the potential. We must solve the KS equations iteratively to a self-consistent solution. The governing mathematical models for the WaveFunction entity thus become the radial Schrödinger and Dirac equations, solved within this self-consistent loop. The discussion thus far does not restrict the number of atoms, however, for the remainder of the chapter we consider single atom systems, though we consider all electrons without approximations typically applied to larger systems 777often treated with non-relativistic methods since core electrons are represented with a pseudo-potential or projected augmented wave [71].
@@ 3.2.2 The Radial Dirac Equation
For heavy atoms, the appropriate single-particle theory requires a relativistic treatment. The central-field Dirac equation leads to two coupled first-order radial equations for the large and small components $(P_{n\kappa},Q_{n\kappa})$ of a four-component spinor, with the relativistic quantum number $\kappa$ encoding $(l,j)$ [72].
@@
The Dirac Hamiltonian possesses a spectrum unbounded from below, a feature that historically plagued basis-set discretizations with “variational collapse” and spurious states [72, 73]. Typical remedies over the past three decades involve shooting methods [74] which require trial solutions for each eigenfunction and convergence parameteres for the solver grid, complicating the need for robust and computationally efficient solutions. Basis set methods [75, 76, 77, 78] solve for all states at once through diagonalization, but struggle with the spurious states of he Dirac Hamiltonian, despite attempts to mitigate these with changes in basis for large and small components [75, 79, 80, 76, 81, 82], Hamiltonian modifications [83, 77, 78, 84, 85], and boundary value constraints [86, 87, 80, 77].
3.3 Robust finite element solvers for isolated atoms
We now consider a methodology which circumvents the spurious states of the Dirac while providing robust and efficient solutions with a high-order finite element basis. The featom code employs solves the governing equations using a finite element basis, implemented in modern modular Fortran. The radial coordinate is discretized into a mesh of finite elements, and within each element, the solution is expanded in a basis of high-order Lagrange polynomials defined on Gauss-Lobatto nodes. This spectral element approach yields exponential convergence with respect to the polynomial order, providing high accuracy with a relatively small number of basis functions. The success of this approach, however, rests on a cascade of intelligent choices in representation at the mathematical, numerical, and software levels.
The finite element expressions and code derived for the squared Dirac Hamiltonian [88, 83] are novel, and these have the same eigenfunctions as the operator, and remains bounded from below while preserving convergence to the Schrödinger limit [83]. Numerical stability arises from the squared operator providing second derivative terms, while known asymptotic forms near the origin allow side-stepping solving for large and small components of the Dirac wavefunction components. We view these concepts and their synthesis through the lens of varying representations below.
@@ 3.3.1 Layer 1: The Mathematical Representation (Squared Hamiltonian)
Direct discretization of the Dirac Hamiltonian operator poses significant difficulties because its energy spectrum remains unbounded from both above and below, leading to spurious, unphysical solutions. To circumvent this, we utilize a different mathematical representation: we solve the eigenvalue problem for the square of the Dirac Hamiltonian, $(H+\mathbb{I}c^{2})^{2}$. This squared operator shares eigenfunctions with the original operator, and its eigenvalues relate simply as the square of the original eigenvalues, $(E+c^{2})^{2}$. Crucially, the squared operator remains bounded from below. This allows the application of standard variational techniques, like the FEM, without generating spurious states.
A key principle here follows from eliminating the need for kinetic balance and other such constraints, and can instead proceed with Galerkin discretization [89, 90] directly in a polynomial basis, the same for both large and small components.
@@ 3.3.4 Layer 4: The Software Representation (Modern, Maintainable Code)
This robust code is supported by a professional software engineering infrastructure and involved introducing the flexible Meson build system alongside the existing Fortran Package Manager (fpm), establishing a comprehensive automated test harness, and refining continuous integration (CI) pipelines. This focus on the software representation ensures correctness through automated validation, lowers the barrier for collaboration, and guarantees long-term maintainability and scientific reproducibility.
@@ 3.4 From KS to HF: conceptually simple, practically subtle in spherical FE
A simple extension towards an HF/KS implementation in the radial FE code follows a “local-in-the-loop, exact-after” strategy 888gh-26 to featom:
@@ 3.5 Performance and accuracy
Figure 10 demonstrates the systematic, reproducible convergence of the solver for uranium in three complementary regimes: domain truncation, \$p\$- and \$h\$-refinement, and the achieved energy precision.
@@
As designed, the featom code is tailored towards being state-of-the-art as a DFT solver for relativistic calculations, which are otherwise treated only through shooting approaches [74], often in the context of expensive quantum chemical finite difference based [97]. Despite this, for systems with a high number of states, featom outperforms shooting method state of the art calculators as well. Here, we appeal to the transitive nature of benchmarks since the state of the art dftatom compares favorably to Desclaux [98], atompp [99], grasp2k [97], MCHF [100], atompaw [101], PEtot [102] and Elk [103], we assert superiority by directly comparing against dftatom. DFT based formulations are always faster, due to the ability to skip the two particle exchange integral. For instance, a B-spline based Hartree-Fock solver [104] reports, for a toy 3-electron model of Uranium (“Lithium like”, 2s(1)), a time of 1.53 seconds without the Breit interaction while we demonstrate a full 29 state Uranium atom:
$1s^{2}2s^{2}2p^{6}3s^{2}3p^{6}3d^{10}4s^{2}4p^{6}4d^{10}4f^{14}5s^{2}5p^{6}5d^{10}5f^{3}6s^{2}6p^{6}6d^{1}7s^{2}.$
solution in 0.36 seconds. Furthermore, we compare the wall time against the state-of-the-art shooting-method code dftatom. For this DFT calculation of uranium converged to an accuracy of $10^{-6}$ Hartree, featom shows a significant speedup for non-relativistic calculations and competitive performance for relativistic ones, validating the effectiveness of the chosen representations.
@@
The benchmark results in Table 1 reflect the distinct scaling characteristics of the FEM used in featom versus the shooting method used in dftatom.
For the Schrödinger equation, featom is significantly faster (28 ms vs 166 ms) because it solves for all eigenvalues in a given angular momentum channel simultaneously via a single diagonalization of a small matrix ($N_{b}\times N_{b}$). In contrast, dftatom must perform iterative radial integrations for each individual electronic state, which accumulates a higher computational cost for high-Z atoms with many electrons.
For the Dirac equation, the performance shifts. In the FEM framework (featom), the basis set size effectively doubles to represent both the large and small spinor components, resulting in matrices of size $2N_{b}\times 2N_{b}$. Since the cost of the dense eigensolver scales cubically with matrix size ($\mathcal{O}(M^{3})$), the computational effort increases by a factor of roughly 8 ($2^{3}$), alongside increased matrix assembly costs. In contrast, the shooting method (dftatom) only sees a linear increase in cost associated with integrating two coupled first-order ODEs rather than one second-order ODE. Consequently, while featom slows down by a factor of $\sim 12$ (28 ms to 360 ms) due to the cubic scaling penalty, dftatom slows down by only $\sim 1.6$ (166 ms to 276 ms), making it slightly faster in the fully relativistic case for this specific benchmark. Despite the associated overhead, the optimized software ensures featom remains competitive in absolute time compared to dftatom and outstrips other solvers.
3.6 Conclusions
This chapter has established the theoretical foundations—specifically the mean-field approximations of Hartree-Fock and Kohn-Sham density functional theory—requisite for interpreting electronic structure calculations throughout this work. While the specific numerical innovations presented here address the challenge of relativistic atomic precision, the broader investigations into reactive systems presented in the remainder of this thesis rely on these fundamental mean-field principles. The calculations for molecular systems in this thesis are at Hartree-Fock level, with spin-unrestricted formalisms for radical species and restricted closed-shell formalisms for stable intermediates.
Within this theoretical framework, we presented a robust finite element formulation for the solution of the radial Schrödinger, Dirac, and Kohn-Sham equations for isolated atoms. The success of the featom solver exemplifies the central thesis that overcoming computational barriers in physics requires a holistic approach to representation. By moving from the Dirac Hamiltonian to its squared operator, we eliminated the spectral pollution of spurious states that has historically plagued basis-set methods. By transforming the dependent variables to account for asymptotic behavior near the nucleus, we resolved the representational conflict between singular Coulombic potentials and smooth polynomial bases, recovering exponential convergence. The practical outcome of these choices is a solver that occupies a unique niche in the current software ecosystem. As demonstrated by the benchmarks, featom provides sub-second wall times for full relativistic DFT calculations of heavy elements like uranium ($Z=92$), outperforming state-of-the-art shooting methods in the relativistic regime while avoiding the complexity of kinetic balance constraints required by B-spline approaches.
This efficiency is not merely a convenience but should translate eventually towards generating better basis sets for high-fidelity reference data needed to train Gaussian Process surrogates discussed in the previous chapter.
Finally, the implementation of featom underscores the critical importance of software accessibility and reproducibility in computational science. For instance, while reviewing the landscape of relativistic solvers, we attempted to evaluate the BERTHA [105] code but were unable to locate source code, finding only Python wrappers dependent on opaque, closed-source binaries999to say nothing of “closed source” code results. In contrast, featom is provided as a modern, modular, open-source library. This ensures that the distinct layers of representation—mathematical, numerical, and algorithmic—remain transparent, verifiable, and adaptable.
However, the converse is also true, not every “conceptually simple” mapping (e.g. Hartree-Fock $\to$ “a potential”) respects the representation. Where the physics demands an operator (HF exchange), we either keep the operator or solve a re-representation problem, e.g. through an optimized effective potential [96, 106, 107].
We will return to this concept in later chapters; when we formulate efficient reaction-path searches and have Gaussian Process accelerators succeed because their internal representations (forces, curvatures, kernels) are chosen to make the numerics stable and the computation scalable.
This serves to clarify that the “correctness” of a scientific result is inextricably linked to the design of the software that produces it. Physical rigor cannot be separated from software architecture. In the subsequent chapter, we will expand this scope from the design of a single solver to the design of the broader computational frameworks required for complex chemical exploration.
@@ 4.1 Bonding analysis
To unambiguously distinguish covalently bonded molecular fragments from transient non-covalent contacts, a robust analysis of the system’s bonding network is essential. This can be approached from two distinct perspectives, namely a heuristic geometric definition or a more rigorous definition based on the system’s electronic structure. We implement in rgpycrumbs both methods, with a pyvista backend 101010inspired by solvis allowing for a flexible and chemically aware analysis of molecular connectivity.
@@ 4.1.1 Geometric method: Covalent cutoff
The simplest and most computationally efficient method for defining a bonding network is based on geometry. In this approach, we define that a bond exists between two atoms, $i$ and $j$, based on their interatomic distance, $d_{ij}$. Specicially, the bond exists when atoms are closer to each other than a scaled sum of their tabulated covalent radii, $r_{i}$ and $r_{j}$. This relationship is governed by the inequality:
@@
Here, $M$ is a dimensionless scaling multiplier 111111typically between $1.1$ and $1.3$ used to adjust the strictness of the criterion. While this method is extremely fast, and widely available due to being the method used by the ASE [112] graphical user interface (GUI); a purely geometric measure for molecules is a significant drawback, as the base unit of calculations are centered on electrons (Chapter 3). Such measures lack “chemical intuition” and can fail in sterically crowded environments where non-bonded atoms are forced into close proximity, leading to the false identification of covalent bonds (as illustrated in Figure 11 B and C.
@@ 4.1.2 Electronic density: Wiberg bond order
Such a measure can be significantly more robust than the geometric approach as it hinges on the calculation of the actual chemical interactions. It can reliably distinguish between genuine covalent bonds, which have significant shared electron density (typically WBO > 0.7), and close non-covalent contacts, which exhibit negligible bond orders. As shown in Figure 11, the WBO between sterically close but non-bonded atoms evaluates to a near zero value, correctly identifying them as belonging to separate molecular fragments. The primary trade-off for this increased accuracy is the higher computational cost associated with performing the underlying electronic structure calculation, which is largely alleviated by using GFN2-xTB semi-empirical calculation [114] 121212this is still not quick enough for extended systems however..
@@
4.2 RMSD projections for path visualization
@@
To visualize the local structure of the potential energy surface, we interpolate the scattered energies onto the RMSD plane. This is achieved using a Radial Basis Function interpolator with a Thin Plate Spline kernel [118], as implemented in SciPy [115].
[…]
where $\phi(r)=r^{2}\log(r)$ is the thin plate spline radial function and $P(\mathbf{x})$ is a low-degree polynomial term included to ensure solvability. The weights $w_{i}$ are determined by solving a linear system that interpolates the energy values $E_{i}$ at the sampled coordinates $(x_{i},y_{i})$. A smoothing parameter $\lambda$ (set to 0.009 in this work) is applied to the diagonal of the interpolation matrix to handle noise in the optimization data and prevent overfitting to high-energy artifacts. This produces the smooth, physically continuous contour maps used to visualize the topography of the energy landscape, shown in section 4.3.3, Figure 13.
@@ 4.3 EON
Most of the calculations in this thesis go through EON. Rather than exessively modifying the SVN copy, a new release was drafted, v2.8.0 141414accompanying documentation : https://eondocs.org. All the methods presented in this thesis are either the direct result of, or stemmed from the landmark modernization of the EON client code. This multi-year development effort, spawning millions of lines of code and documentation, overhauled the entire framework to be more powerful, flexible, and robust, transforming it from a legacy tool into a modern scientific platform. The effort focused on several key areas of software engineering.
@@ 4.3.1 Eliminating I/O Bottlenecks with a Client-Server Architecture
This effort required not only developing the server architecture within EON 151515gh-244 to EON but also contributing directly to the NWChem codebase to improve its capabilities as a client. A key pull request, which was merged into the official NWChem repository 161616gh-1145 to NWChem , enhanced its socket client with a robust polling and retry mechanism. This modification allows the NWChem client to wait patiently for the EON server to become available, a critical feature for ensuring stable, loosely-coupled communication between the two persistent programs.